kathy earns $445 a week for 40 hours of work and $25 and hour for each hour over 40. how many hours did kathy work if she earned $570 in one week?
Answer by checkley77(12844)

445+25X=570
25X=570-445
25X=125
X=125/25
X=5 HOURS WORKED OVERTIME.